# Distributed MIMO Positioning: Fundamental Limit Analysis and User Tracking Framework Design

### v1

@@ III-D Numerical Results
This operation normalizes the energy of each element to one while preserving the intra-element variations across frequencies, polarizations, and directions, as well as the inter-element amplitude differences. Unless otherwise specified, we assume a single vertically polarized UE in the simulations with antenna responses $c_{R_{V}}=1$ and $c_{R_{H}}=0$, and a tilt angle of $\beta=45^{\circ}$. The channel amplitude $\alpha_{VV}$ is modeled based on the free-space path loss, given by $\alpha_{VV}=\frac{\lambda}{4\pi\|\mathbf{P}^{\text{AP}}-\mathbf{P}^{\text{UE}}\|}$. The transmission signal power is set to $E_{s}=\frac{|\alpha_{VV}|^{2}}{NN_{f}}$. We define the signal-to-noise ratio (SNR) as $\text{SNR}=\frac{E_{s}}{\sigma^{2}}$.
@@ III-D1 Impact of Array Aperture
We consider a single AP located at $\mathbf{r}^{\text{AP}}=[0,0,0]^{\text{T}}$ with rotation $\mathbf{r}^{\text{AP}}=\mathbf{I}_{3\times 3}$. The UE is fixed at $\mathbf{p}^{\text{UE}}=[2,2,2]^{\text{T}}.$ Here, the root mean squared error (RMSE) of the estimated UE position, achieved by a maximum likelihood estimator, is also evaluated compared to the derived PEB. The RMSE is defined as $\text{RMSE}(\mathbf{p}^{\text{UE}})=\mathbb{E}\left\{\sqrt{\|\hat{\mathbf{p}}^{\text{UE}}-\mathbf{p}^{\text{UE}}\|^{2}}\right\}$, where $\hat{\mathbf{p}}^{\text{UE}}$ is the estimated UE position. The numerical RMSE is obtained from 1000 Monte Carlo experiments.
@@ III-D3 Impact of Spatial Distribution of APs
To investigate the relationship between the PEBs and the spatial distributions of APs, we first analyze the PEB with respect to the defined AP geometry factor $D\left(\mathbf{p}^{\text{AP}}_{1},...,\mathbf{p}^{\text{AP}}_{K}\right)$ in Theorem 3. A 2D scenario with four distributed APs, each equipped with the described sub-6 GHz UPA, is considered. The AP positions are randomly generated, while the propagation distances between each AP and the UE are fixed at 5 m under different SNRs, i.e., 10, 15, 20, and 25 dB. The angle between each AP and the UE is assumed to remain unchanged. Fig. 5 shows the relationships between the associated PEB and $D\left(\mathbf{p}^{\text{AP}}_{1},...,\mathbf{p}^{\text{AP}}_{4}\right)$. It can be observed that the PEB exhibits a non-decreasing trend with $D\left(\mathbf{p}^{\text{AP}}_{1},...,\mathbf{p}^{\text{AP}}_{4}\right)$. When $\epsilon=\frac{[\mathbf{F}_{\Theta_{k}}]_{\theta\theta}}{\tau^{2}[\mathbf{F}_{\Theta_{k}}]_{\tau\tau}}=1$, the distance and azimuth information intensities of each AP are equal. In this case, the information ellipse becomes circular, as discussed earlier, indicating that the geometric distribution of the APs does not affect the limits of positioning performance. These results validate that the defined factor $D\left(\mathbf{p}^{\text{AP}}_{1},...,\mathbf{p}^{\text{AP}}_{K}\right)$ can serve as a sufficient indicator to evaluate the effect of the spatial distribution of APs on the PEB.
@@ IV-A1 State Model
The state of the $k$-th AP includes its 3D location $\mathbf{p}^{\text{AP}}_{k}$ and 3D rotation $\mathbf{r}_{k}^{\text{AP}}$, which are considered a priori. The state $\mathbf{x}_{m}^{t}$ of the $m$-th mobile UE at time step $t$ is characterized by its 3D position $\mathbf{p}^{\text{UE}}_{m}(t)$ and velocity $\mathbf{\upsilon}_{m}^{t}$. The state transition density is expressed as
@@ IV-A2 Measurement Model
To solve the UE tracking problem, we consider a two-stage process. First, after receiving the signals, channel estimation is performed at each AP. Second, the estimated MPC delay and AoA are obtained for subsequent UE tracking. Assuming that the measurements of different APs are independent and homogeneous, the delays and AoAs are transformed into position-related measurements $\mathbf{z}$ in the state space. This is achieved using the inverse sigma points [38], which is common practice to linearize non-linear measurements. Please refer to Algorithm 1 in [1] for more details. The resulting measurement model for the UE tracking is expressed as $\mathbf{z}_{l,k,m}^{t}=\mathbf{p}^{\text{UE}}_{m}(t)+\mathbf{n}^{t}_{l,k,m}$, where $\mathbf{n}^{t}_{l,k,m}$ denotes the measurement noise with a covariance of $\mathbf{W}^{t}_{l,k,m}$. The likelihood function is thus expressed as
$h\left(\mathbf{z}_{l,k,m}^{t}|\mathbf{x}_{m}^{t}\right)=\mathcal{N}\left(\mathbf{z}_{l,k,m}^{t};\xi\left(\mathbf{x}_{m}^{t}\right),\mathbf{W}^{t}_{l,k,m}\right),$ (36)
with $\xi\left(\mathbf{x}_{m}^{t}\right)=\mathbf{p}^{\text{UE}}_{m}(t)$. Note that there are non-LoS (NLoS) paths in the received measurements, i.e., MPCs that may have experienced reflection, scattering, or diffraction. In this work, such paths are treated as clutter [^1] While there are methods that utilize NLoS paths for UE tracking, as discussed in [23, 24, 25, 26], such approaches are beyond the scope of this work.. Furthermore, it is unclear which measurements originate directly from the UEs or from clutter, making it difficult to perform an accurate data association. To manage this uncertainty, all measurements received by the $k$-th AP at time $t$ are modeled as an RFS, i.e., $\mathbf{Z}_{k}^{t}=\left\{\mathbf{z}_{1,k}^{t},...,\mathbf{z}_{L,k}^{t}\right\}$, $L=\sum_{m}^{M}\hat{L}_{k,m}$. The set of UE states at time $t$ is modeled similarly to another RFS $\mathbf{X}^{t}=\left\{\mathbf{x}_{1}^{t},...,\mathbf{x}_{M}^{t}\right\}$. Note that $\hat{L}_{k,m}$ can differ from the true number of MPCs $L_{k,m}$ in (2) due to missed detections or false alarms. Since LoS paths may also be undetected in some time steps, a detection probability $p_{\text{D}}\in[0,1]$ is introduced to quantify the probability that a given measurement comes directly from a UE. Based on this model, we employ a PHD filter, where the RFS $\mathbf{Z}_{k}^{t}$ is used to estimate the UE state RFS $\mathbf{X}^{t}$, as will be detailed in the following section.
@@ IV-B Global PHD Filter for UE Tracking
Based on a Bayesian filtering framework, a PHD filter aims to recursively estimate the first-order statistical moment of the posterior density of the RFS $p(\mathbf{X}^{t}|\mathbf{Z}^{t}_{k})$, referred to as the PHD $v_{t}(\mathbf{x})$. The local maxima of $v_{t}(\mathbf{x})$ provide estimates of the RFS $\mathbf{X}^{t}$ of the UE state, allowing for the extraction of individual UE states. Specifically, the PHD propagates over time through the following prediction and update steps:
@@
$\begin{split}v_{t}(\mathbf{x})=&(1-p_{D})v_{t|t-1}(\mathbf{x})\\ &+\sum_{\mathbf{z}\in\mathbf{Z}^{t}_{k}}\frac{p_{D}h_{t}(\mathbf{z}|\xi(\mathbf{x}))v_{t|t-1}(\mathbf{x})}{\lambda_{c}+p_{D}\int h_{t}(\mathbf{z}|\xi(\mathbf{x}))v_{t|t-1}(\mathbf{x})\text{d}\mathbf{x}},\end{split}$ (38)
@@
A key feature of the introduced filter lies in its multi-source measurement transformation before each PHD update. That is, AoA and delay measurements from different APs are first transformed into position-domain measurements in the same state space. Then we adopt the clustering method in [39] to group those multi-source position-domain measurements corresponding to the same objects and generate a ‘proxy’ measurement to be used in the PHD update. This differs from existing filters [23, 24, 25], in which AoA and delay measurements are used directly in the nonlinear measurement model, leading to the need to approximate them through a nonlinear filter, which generally exhibits high computational complexity.
@@ IV-C PEB-Aware AP Management
An effective strategy to determine which APs are active at each time step is essential to maintain continuous UE tracking while minimizing system overhead. Existing antenna selection strategies for UE tracking typically fall into two categories: FoV-based and error bound-based selection. The FoV-based approach dynamically selects APs that can cover current UE locations [1]. This method is computationally efficient; however, it faces practical limitations, especially when APs are equipped with realistic antenna arrays, which makes accurate FoV modeling difficult. The error bound-based approach selects APs that are expected to provide a lower error bound for tracking. In this context, performance bounds such as the posterior CRLB [28] and the Bayesian CRLB [29] are widely used as management criteria. However, these bounds cannot be directly computed from the PHD filter due to the absence of explicit data association between measurements and targets or clutter. In this work, we use the PEB derived in Section III. Since the PEB reflects how accurately a UE can be localized given its relative geometry to the APs, it inherently indicates the maximum achievable quality of the measurements. Therefore, APs associated with lower PEBs are more likely to contribute to high-quality measurements and ultimately to improved tracking performance.
@@ Appendix A Derivation of the Elements of $\mathbf{F}_{\Theta_{k}}$
$\frac{\partial\mathbf{D}}{\partial\theta}=\sum_{\text{p}\in\{V,H\}}\alpha_{\text{pp}}\cdot c_{T_{\text{p}}}^{\prime}\left[\left(\mathbf{Q}_{R_{\text{p}}}\cdot\left[\left(j\mathbf{\alpha}_{\theta}\odot\mathbf{w}_{\theta}\right)\otimes\mathbf{w}_{\phi}\right]\right)\otimes\mathbf{b}_{f}\right],$
$\frac{\partial\mathbf{D}}{\partial\phi}=\sum_{\text{p}\in\{V,H\}}\alpha_{\text{pp}}\cdot c_{T_{\text{p}}}^{\prime}\left[\left(\mathbf{Q}_{R_{\text{p}}}\cdot\left[\mathbf{w}_{\theta}\otimes\left(j\mathbf{\alpha}_{\phi}\odot\mathbf{w}_{\phi}\right)\right]\right)\otimes\mathbf{b}_{f}\right],$
$\frac{\partial\mathbf{D}}{\partial\tau}=\sum_{\text{p}\in\{V,H\}}\alpha_{\text{pp}}\cdot c_{T_{\text{p}}}^{\prime}\left[\left(\mathbf{Q}_{R_{\text{p}}}\cdot\left[\mathbf{w}_{\theta}\otimes\mathbf{w}_{\phi}\right]\right)\otimes(-j2\pi\mathbf{f}\odot\mathbf{b}_{f})\right],$
$\frac{\partial\mathbf{D}}{\partial\mathcal{R}\{\alpha_{VV/HH}\}}=c^{\prime}_{T_{\text{V/H}}}\left(\mathbf{b}_{t}\otimes[\mathbf{Q}_{R_{\text{V/H}}}\cdot(\mathbf{w}_{\theta}\otimes\mathbf{w}_{\phi})]\otimes\mathbf{b}_{f}\right),$
$\frac{\partial\mathbf{D}}{\partial\mathcal{I}\{\alpha_{VV/HH}\}}=jc^{\prime}_{T_{\text{V/H}}}\left(\mathbf{b}_{t}\otimes[\mathbf{Q}_{R_{\text{V/H}}}\cdot(\mathbf{w}_{\theta}\otimes\mathbf{w}_{\phi})]\otimes\mathbf{b}_{f}\right),$
@@ Appendix D Proof of Theorem 4
$\begin{split}\left[\mathbf{F}_{\Theta_{k}}\right]_{\theta\theta}=&\frac{2E_{s}}{\sigma^{2}}\mathcal{R}\big\{\cos^{2}\beta(\mathbf{A}^{\theta}_{V})^{\text{H}}\mathbf{A}^{\theta}_{V}+\sin^{2}\beta(\mathbf{A}^{\theta}_{H})^{\text{H}}\mathbf{A}^{\theta}_{H}\\ &+\cos\beta\sin\beta\left((\mathbf{A}^{\theta}_{H})^{\text{H}}\mathbf{A}^{\theta}_{V}+(\mathbf{A}^{\theta}_{V})^{\text{H}}\mathbf{A}^{\theta}_{H}\right)\big\},\end{split}$ (62)
where $(\mathbf{A}^{\theta}_{\text{p}})^{\text{H}}\mathbf{A}^{\theta}_{\text{p}^{\prime}}=\alpha_{\text{pp}}^{*}\alpha_{\text{p}^{\prime}\text{p}^{\prime}}\big(\left[(\mathbf{\alpha}_{\theta}\odot\mathbf{w}_{\theta})\otimes\mathbf{w}_{\phi}\right]\cdot\mathbf{Q}_{T_{\text{p}}}^{\text{H}}\mathbf{Q}_{T_{\text{p}^{\prime}}}\cdot\left[(\mathbf{\alpha}_{\theta}\odot\mathbf{w}_{\theta})\otimes\mathbf{w}_{\phi}\right]^{\text{H}}\big).$ When $\left(\mathbf{c}_{V}\right)^{\mathrm{H}}\mathbf{c}_{H}+\left(\mathbf{c}_{H}\right)^{\mathrm{H}}\mathbf{c}_{V}\ll\left(\mathbf{c}_{V}\right)^{\mathrm{H}}\mathbf{c}_{V}+\left(\mathbf{c}_{H}\right)^{\mathrm{H}}\mathbf{c}_{H}$, we have $(\mathbf{A}^{\theta}_{H})^{\text{H}}\mathbf{A}^{\theta}_{V}+(\mathbf{A}^{\theta}_{V})^{\text{H}}\mathbf{A}^{\theta}_{H}\ll(\mathbf{A}^{\theta}_{H})^{\text{H}}\mathbf{A}^{\theta}_{H}+(\mathbf{A}^{\theta}_{V})^{\text{H}}\mathbf{A}^{\theta}_{V}.$ Therefore, $\mathbb{E}\{\left[\mathbf{F}_{\Theta_{k}}\right]_{\theta\theta}\}\approx\frac{2E_{s}}{\sigma^{2}}\mathcal{R}\bigg\{\mathbb{E}\{\cos^{2}\beta\}(\mathbf{A}^{\theta}_{V})^{\text{H}}\mathbf{A}^{\theta}_{V}+\mathbb{E}\{\sin^{2}\beta\}(\mathbf{A}^{\theta}_{H})^{\text{H}}\mathbf{A}^{\theta}_{H}\bigg\}.$ Since $\beta\sim\mathcal{VM}(\mu,\kappa)$, there is $\mathbb{E}[\cos(n\beta)]=\frac{I_{n}(k)}{I_{0}(k)}\cos(n\mu)$, thus we have

### v2

@@ III-D Numerical Results
This operation normalizes the expected energy of each element to one while preserving the intra-element variations across frequencies, polarizations, and directions, as well as the inter-element amplitude differences. Unless otherwise specified, we assume a single vertically polarized UE in the simulations with antenna responses $c_{\rm{R_{V}}}=1$ and $c_{\rm{R_{H}}}=0$, and a tilt angle of $\beta=45^{\circ}$. The channel amplitude $\alpha_{\rm{VV}}$ is modeled based on the free-space path loss, given by $\alpha_{\rm{VV}}=\frac{\lambda}{4\pi\|\mathbf{P}^{\text{AP}}-\mathbf{P}^{\text{UE}}\|}$. The transmission signal power is set to $E_{s}=\frac{|\alpha_{\rm{VV}}|^{2}}{NN_{f}}$. We define the receive signal-to-noise ratio (SNR) as $\frac{E_{s}}{\sigma^{2}}$.
@@ III-D1 Impact of Array Aperture
We consider a single AP located at $\mathbf{p}^{\text{AP}}=[0,0,0]^{\text{T}}$ with orientation $\mathbf{r}^{\text{AP}}=\mathbf{I}_{3\times 3}$. The UE is fixed at $\mathbf{p}^{\text{UE}}=[2,2,2]^{\text{T}}.$ Here, the root mean squared error (RMSE) of the estimated UE position, achieved by a maximum likelihood estimator, is also evaluated compared to the derived PEB. The RMSE is defined as $\text{RMSE}(\mathbf{p}^{\text{UE}})=\mathbb{E}\left\{\sqrt{\|\hat{\mathbf{p}}^{\text{UE}}-\mathbf{p}^{\text{UE}}\|^{2}}\right\}$, where $\hat{\mathbf{p}}^{\text{UE}}$ is the estimated UE position. The numerical RMSE is obtained from 1000 Monte Carlo experiments.
@@ III-D3 Impact of Spatial Distribution of APs
To investigate the relationship between the PEBs and the spatial distributions of APs, we first analyze the PEB with respect to the defined AP geometry factor $D\left(\mathbf{p}^{\text{AP}}_{1},...,\mathbf{p}^{\text{AP}}_{K}\right)$ in Theorem 3. A 2D scenario with four distributed APs, each equipped with the described sub-6 GHz UPA, is considered. The AP positions are randomly generated, while the propagation distances between each AP and the UE are fixed at 5 m under different receive SNRs, i.e., 10, 15, 20, and 25 dB. The angle between each AP and the UE is assumed to remain unchanged. Fig. 5 shows the relationships between the associated PEB and $D\left(\mathbf{p}^{\text{AP}}_{1},...,\mathbf{p}^{\text{AP}}_{4}\right)$. It can be observed that the PEB exhibits a non-decreasing trend with $D\left(\mathbf{p}^{\text{AP}}_{1},...,\mathbf{p}^{\text{AP}}_{4}\right)$. When $\epsilon=\frac{[\mathbf{F}_{\Theta_{k}}]_{\theta\theta}}{\tau^{2}[\mathbf{F}_{\Theta_{k}}]_{\tau\tau}}=1$, the distance and azimuth information intensities of each AP are equal. In this case, the information ellipse becomes circular, as discussed earlier, indicating that the geometric distribution of the APs does not affect the limits of positioning performance. These results validate that the defined factor $D\left(\mathbf{p}^{\text{AP}}_{1},...,\mathbf{p}^{\text{AP}}_{K}\right)$ can serve as a sufficient indicator to evaluate the effect of the spatial distribution of APs on the PEB.
@@ IV-A1 State Model
The state of the $k$-th AP includes its 3D location $\mathbf{p}^{\text{AP}}_{k}$ and 3D orientation $\mathbf{r}_{k}^{\text{AP}}$, which are considered a priori. The state $\mathbf{x}_{m}^{t}$ of the $m$-th mobile UE at time step $t$ is characterized by its 3D position $\mathbf{p}^{\text{UE}}_{m}(t)$ and velocity $\mathbf{\upsilon}_{m}^{t}$. The state transition density is expressed as
@@ IV-A2 Measurement Model
To solve the UE tracking problem, we consider a two-stage process. First, after receiving the signals, channel estimation is performed at each AP, and the estimated MPC delays and AoAs are obtained. Second, assuming that the measurements of different APs are independent and homogeneous, the delays and AoAs are transformed into position-related measurements $\tilde{\mathbf{Z}}^{t}=\{\tilde{\mathbf{z}}_{1,k}^{t},...,\tilde{\mathbf{z}}_{L,k}^{t}\}_{k=1}^{K}$. This is achieved using the inverse sigma points [38], which is a common practice to linearize non-linear measurements. Then the clustering method in [39] is adopted to group those multi-source position-related measurements corresponding to the same objects and generate a ‘proxy’ measurement. Let $\mathbf{Z}^{t}=\{\mathbf{z}_{1}^{t},...,\mathbf{z}_{L^{\prime}}^{t}\}$ denote the set of these ‘proxy’ measurements which are used in the PHD update.
The measurement model for the UE tracking is expressed as $\mathbf{z}_{l,m}^{t}=\mathbf{p}^{\text{UE}}_{m}(t)+\mathbf{n}^{t}_{l,m}$, where $\mathbf{n}^{t}_{l,m}$ denotes the measurement noise with a covariance of $\mathbf{W}^{t}_{l,m}$. The likelihood function is thus expressed as
$h\left(\mathbf{z}_{l,m}^{t}|\mathbf{x}_{m}^{t}\right)=\mathcal{N}\left(\mathbf{z}_{l,m}^{t};\xi\left(\mathbf{x}_{m}^{t}\right),\mathbf{W}^{t}_{l,m}\right),$ (36)
with $\xi\left(\mathbf{x}_{m}^{t}\right)=\mathbf{p}^{\text{UE}}_{m}(t)$. Note that there are non-LoS (NLoS) paths in the received measurements, i.e., MPCs that may have experienced reflection, scattering, or diffraction. In this work, such paths are treated as clutter [^1] While there are methods that utilize NLoS paths for UE tracking, as discussed in [23, 24, 25, 26], such approaches are beyond the scope of this work.. Furthermore, it is unclear which measurements originate directly from the UEs or from clutter, making it difficult to perform an accurate data association. To manage this uncertainty, the set of UE states at time $t$ is modeled as an RFS $\mathbf{X}^{t}=\left\{\mathbf{x}_{1}^{t},...,\mathbf{x}_{M}^{t}\right\}$. Since LoS paths may also be undetected in some time steps, a detection probability $p_{\text{D}}\in[0,1]$ is introduced to quantify the probability that a given measurement comes directly from a UE. Based on this model, we employ a PHD filter, where the measurement sets $\mathbf{Z}^{t}$ are used to estimate the UE state RFS $\mathbf{X}^{t}$, as will be detailed in the following section.
@@ IV-B Global PHD Filter for UE Tracking
Based on a Bayesian filtering framework, a PHD filter aims to recursively estimate the first-order statistical moment of the posterior density of the RFS $p(\mathbf{X}^{t}|\mathbf{Z}^{t})$, referred to as the PHD $v_{t}(\mathbf{x})$. The local maxima of $v_{t}(\mathbf{x})$ provide estimates of the RFS $\mathbf{X}^{t}$ of the UE state, allowing for the extraction of individual UE states. Specifically, the PHD propagates over time through the following prediction and update steps:
@@
$\begin{split}v_{t}(\mathbf{x})=&(1-p_{D})v_{t|t-1}(\mathbf{x})\\ &+\sum_{\mathbf{z}\in\mathbf{Z}^{t}}\frac{p_{D}h_{t}(\mathbf{z}|\xi(\mathbf{x}))v_{t|t-1}(\mathbf{x})}{\lambda_{c}+p_{D}\int h_{t}(\mathbf{z}|\xi(\mathbf{x}))v_{t|t-1}(\mathbf{x})\text{d}\mathbf{x}},\end{split}$ (38)
@@
We recall that a key feature of the introduced global filter lies in its multi-source measurement transformation before each PHD update. That is, AoA and delay measurements from different APs are first transformed into position-related measurements. Then, those multi-source position-domain measurements corresponding to the same objects are represented by a ‘proxy’ measurement. This differs from existing filters [23, 24, 25], in which AoA and delay measurements are used directly in the nonlinear measurement model, leading to the need to approximate them through a nonlinear filter, which generally exhibits high computational complexity.
@@ IV-C PEB-Aware AP Management
An effective strategy to determine which APs are active at each time step is essential to maintain continuous UE tracking while minimizing system overhead. Existing antenna selection strategies for UE tracking typically fall into two categories: FoV-based and error bound-based selection. The FoV-based approach dynamically selects APs that can cover current UE locations [1]. This method is computationally efficient; however, it faces practical limitations, especially when APs are equipped with realistic antenna arrays, which makes accurate FoV modeling difficult. The error bound-based approach selects APs that are expected to provide a lower error bound for tracking. In this context, performance bounds such as the posterior CRLB [28] and the Bayesian CRLB [29] are widely used as management criteria. However, these bounds cannot be directly computed if the explicit data association between measurements and targets or clutter is unknown, which is a key feature of the PHD filter. In this work, we use the PEB derived in Section III. Since the PEB reflects how accurately a UE can be localized given its relative geometry to the APs, it inherently indicates the maximum achievable quality of the measurements. Therefore, APs associated with lower PEBs are more likely to contribute to high-quality measurements and ultimately to improved tracking performance.
@@ Appendix A Derivation of the Elements of $\mathbf{F}_{\Theta_{k}}$
$\frac{\partial\mathbf{D}}{\partial\theta}=\sum_{\text{p}\in\{V,H\}}\alpha_{\text{pp}}\cdot c_{\rm{T}_{\text{p}}}^{\prime}\left[\left(\mathbf{Q}_{\rm{R}_{\text{p}}}\cdot\left[\left(j\mathbf{\alpha}_{\theta}\odot\mathbf{w}_{\theta}\right)\otimes\mathbf{w}_{\phi}\right]\right)\otimes\mathbf{b}_{f}\right],$
$\frac{\partial\mathbf{D}}{\partial\phi}=\sum_{\text{p}\in\{V,H\}}\alpha_{\text{pp}}\cdot c_{\rm{T}_{\text{p}}}^{\prime}\left[\left(\mathbf{Q}_{\rm{R}_{\text{p}}}\cdot\left[\mathbf{w}_{\theta}\otimes\left(j\mathbf{\alpha}_{\phi}\odot\mathbf{w}_{\phi}\right)\right]\right)\otimes\mathbf{b}_{f}\right],$
$\frac{\partial\mathbf{D}}{\partial\tau}=\sum_{\text{p}\in\{V,H\}}\alpha_{\text{pp}}\cdot c_{\rm{T}_{\text{p}}}^{\prime}\left[\left(\mathbf{Q}_{\rm{R}_{\text{p}}}\cdot\left[\mathbf{w}_{\theta}\otimes\mathbf{w}_{\phi}\right]\right)\otimes(-j2\pi\mathbf{f}\odot\mathbf{b}_{f})\right],$
$\frac{\partial\mathbf{D}}{\partial\mathcal{R}\{\alpha_{\rm{VV/HH}}\}}=c^{\prime}_{\rm{T}_{\text{V/H}}}\left(\mathbf{b}_{t}\otimes[\mathbf{Q}_{\rm{R}_{\text{V/H}}}\cdot(\mathbf{w}_{\theta}\otimes\mathbf{w}_{\phi})]\otimes\mathbf{b}_{f}\right),$
$\frac{\partial\mathbf{D}}{\partial\mathcal{I}\{\alpha_{\rm{VV/HH}}\}}=jc^{\prime}_{\rm{T}_{\text{V/H}}}\left(\mathbf{b}_{t}\otimes[\mathbf{Q}_{\rm{R}_{\text{V/H}}}\cdot(\mathbf{w}_{\theta}\otimes\mathbf{w}_{\phi})]\otimes\mathbf{b}_{f}\right),$
@@ Appendix D Proof of Theorem 4
$\begin{split}\left[\mathbf{F}_{\Theta_{k}}\right]_{\theta\theta}=&\frac{2E_{s}}{\sigma^{2}}\mathcal{R}\big\{\cos^{2}\beta(\mathbf{A}^{\theta}_{\rm{V}})^{\text{H}}\mathbf{A}^{\theta}_{\rm{V}}+\sin^{2}\beta(\mathbf{A}^{\theta}_{\rm{H}})^{\text{H}}\mathbf{A}^{\theta}_{\rm{H}}\\ &+\cos\beta\sin\beta\left((\mathbf{A}^{\theta}_{\rm{H}})^{\text{H}}\mathbf{A}^{\theta}_{\rm{V}}+(\mathbf{A}^{\theta}_{\rm{V}})^{\text{H}}\mathbf{A}^{\theta}_{\rm{H}}\right)\big\},\end{split}$ (62)
where $(\mathbf{A}^{\theta}_{\text{p}})^{\text{H}}\mathbf{A}^{\theta}_{\text{p}^{\prime}}=\alpha_{\text{pp}}^{*}\alpha_{\text{p}^{\prime}\text{p}^{\prime}}\big(\left[(\mathbf{\alpha}_{\theta}\odot\mathbf{w}_{\theta})\otimes\mathbf{w}_{\phi}\right]\cdot\mathbf{Q}_{T_{\text{p}}}^{\text{H}}\mathbf{Q}_{T_{\text{p}^{\prime}}}\cdot\left[(\mathbf{\alpha}_{\theta}\odot\mathbf{w}_{\theta})\otimes\mathbf{w}_{\phi}\right]^{\text{H}}\big).$ When $\left(\mathbf{c}_{\rm{V}}\right)^{\mathrm{H}}\mathbf{c}_{\rm{H}}+\left(\mathbf{c}_{\rm{H}}\right)^{\mathrm{H}}\mathbf{c}_{\rm{V}}\ll\left(\mathbf{c}_{\rm{V}}\right)^{\mathrm{H}}\mathbf{c}_{\rm{V}}+\left(\mathbf{c}_{\rm{H}}\right)^{\mathrm{H}}\mathbf{c}_{\rm{H}}$, we have $(\mathbf{A}^{\theta}_{\rm{H}})^{\text{H}}\mathbf{A}^{\theta}_{\rm{V}}+(\mathbf{A}^{\theta}_{\rm{V}})^{\text{H}}\mathbf{A}^{\theta}_{\rm{H}}\ll(\mathbf{A}^{\theta}_{\rm{H}})^{\text{H}}\mathbf{A}^{\theta}_{\rm{H}}+(\mathbf{A}^{\theta}_{\rm{V}})^{\text{H}}\mathbf{A}^{\theta}_{\rm{V}}.$ Therefore, $\mathbb{E}\{\left[\mathbf{F}_{\Theta_{k}}\right]_{\theta\theta}\}\approx\frac{2E_{s}}{\sigma^{2}}\mathcal{R}\bigg\{\mathbb{E}\{\cos^{2}\beta\}(\mathbf{A}^{\theta}_{\rm{V}})^{\text{H}}\mathbf{A}^{\theta}_{\rm{V}}+\mathbb{E}\{\sin^{2}\beta\}(\mathbf{A}^{\theta}_{\rm{H}})^{\text{H}}\mathbf{A}^{\theta}_{\rm{H}}\bigg\}.$ Since $\beta\sim\mathcal{VM}(\mu,\kappa)$, there is $\mathbb{E}[\cos(n\beta)]=\frac{I_{n}(k)}{I_{0}(k)}\cos(n\mu)$, thus we have